OpenStudy (anonymous):

GAIN A MEDAL! write an equation for each vertical translation of y= f(x) 2/3 units down 4 units up 2 units up

jigglypuff314 (jigglypuff314):

vertical translation means to move up and down so for "2/3 units down" the function would be y = x - 2/3 y = x being the parent function and minus 2/3 because going down

OpenStudy (anonymous):

y=f(x)-2/3 y=f(x)+4 y=f(x)+2